OpenStudy (anonymous):

Solve for x: 1/x+3 - 2/3-x = 4/x^2-9

OpenStudy (anonymous):

As in 1 OVER x+3 and 2 OVER 3-x

OpenStudy (anonymous):

is it 4/(x^2-9)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay then x = 1/3

OpenStudy (anonymous):

How'd you do that? o_o

OpenStudy (anonymous):

well first thing you want to do is try and get a common factor in the denominator

OpenStudy (anonymous):

so what we do is subtract 4/(x^2-9) to the left side

OpenStudy (anonymous):

Ohh! I forgot about that part

OpenStudy (anonymous):

so we now have: \[1/(x+3) - 2/(-x+3) - 4/(x^2-9)\] So we multiply the first term by (x+3)/(x+3) and the second term by (-x+3)/(-x+3) This gives us \[[(x+3)+(2x-6)-4] / (x^2-9) = 0\]

OpenStudy (anonymous):

then simplifies down to (3x-1)/ (x^2-9) = 0 Multiply both sides by (x^2-9) so your left with 3x-1=0 and solve for x so x = 1/3

OpenStudy (anonymous):

Long but I hope it helped

OpenStudy (anonymous):

thank you again. I'd bake you cookies if I could.

OpenStudy (anonymous):

haha thanks i appreciate it and do love cookies lol